OpenStudy (anonymous):

8. A company manufactures and sells video games. A survey of video game stores indicated that at a price of $71 each, the demand would be 200 games, and at a price of $31 each, the demand would be 600 games. If a linear relationship between price and demand exists, which of the following equations models the price-demand relationship? (Let x represent the price per video game and y represent the demand.) @Coolsector all I need is that formula you gave me earlier

OpenStudy (anonymous):

we have two "points" (71,200) , (31,600) so the slope m is given by (600-200) / (31-71) = -10 using y=y1+m(x-x1) we get y = 200 -10x +710 y = -10x + 910

OpenStudy (anonymous):

Hmph I wanted to try it for myself. :p But thanks for the answer :D

OpenStudy (anonymous):

ok im sorry didnt know what you really need

OpenStudy (anonymous):

Oh well thanks anyhow :)